CRYSTAL FIELD THEORY ANALYSIS OF ROVIBRATIONAL SPECTRA OF CARBON MONOXIDE MONOMERS IN SOLID PARAHYDROGEN

\underline{MARIO E. FAJARDO}, AFRL/MNME, Energetic Materials Branch, Ordnance Division, U.S. Air Force Research Lab, 2306 Perimeter Rd., Eglin AFB, FL 32542-5910. mario.fajardo@eglin.af.mil; TAKAMASA MOMOSE, Division of Chemistry, Graduate School of Science, Kyoto University, Kyoto, 605-8502, Japan.

We report our analysis of high-resolution rovibrational infrared absorption spectra of carbon monoxide (CO) molecules isolated in solid parahydrogen (pH_{2}). Experimental data, comprising roughly a dozen transitions for each isotopomer, are presented for ^{12}C^{16}O, ^{13}C^{16}O, ^{12}C^{18}O, and ^{13}C^{18}O. Assignments are made using a Hamiltonian derived using crystal field theory for a diatomic rotor in a D_{3h} symmetry single-substitutional trapping site in hexagonal-close-packed solid pH_{2}. Successful fitting of these spectra requires treating the CO rotational constants (B_{e}, D_{e}, α_{e}) as adjustable parameters; the resulting B_{e}(pH_{2}) values are reduced to roughly 80% of their gas phase values.